# Rethinking Irregular Time Series Forecasting: A Simple yet Effective Baseline

### v3

@@ 3.4 Query-based Weighted Aggregation
$s_{p}^{n}$ $=\frac{q^{n}\cdot(h_{pe,p}^{n})^{T}}{\sqrt{D}}\quad\text{for }p=1,\dots,P,$ (6)
$\beta^{n}$ $=\text{Softmax}(\{s_{p}^{n}\}_{p=1}^{P}),$ (7)
$H_{c}^{n}$ $=\sum_{p=1}^{P}\beta_{p}^{n}h_{pe,p}^{n},$ (8)
The resulting summary representations for all channels, $H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$. A layer normalization is then applied to stabilize the input for the decoder. This lightweight aggregation mechanism is a direct testament to our design philosophy: by front-loading the complexity into TAPA, the subsequent modules can be elegantly simple yet powerful.
@@ 3.5 Forecasting Decoder
The final step of the APN framework is the ultimate demonstration of its architectural elegance. Having distilled the entire irregular history into a powerful, fixed-size representation $H_{final}^{n}$, the forecasting task becomes remarkably straightforward.
The prediction for a query time $\tau_{k}$ is made by a simple two-layer MLP decoder. This decoder takes the concatenated summary representation $H_{final}^{n}$ and the learnable temporal encoding of the query time, $TE(\tau_{k})$, as input to produce the final value $\hat{v}^{n}_{k}$:
$\hat{v}^{n}_{k}=\text{MLP}(\text{Concat}(H_{final}^{n},TE(\tau_{k})))\in\mathbb{R}$ (9)
@@
$\mathcal{L}=\frac{1}{N\cdot Q}\sum_{n=1}^{N}\sum_{k=1}^{Q}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (10)
where $Q$ is the number of query points for each channel.
@@ 4.1 Setup
Datasets: To comprehensively evaluate the model’s performance, we select four widely-used IMTS datasets, including PhysioNet, MIMIC, Activity, and USHCN. These datasets span multiple domains such as healthcare, biomechanics, and climate science. More details of the benchmark datasets are included in Appendix A.1.
Baselines: To comprehensively evaluate the performance of APN, we compare it with nineteen baseline models. These baseline models can be broadly categorized into the following groups: (1) Regular MTSF models: including DLinear (Zeng et al. 2023), TimesNet (Wu et al. 2023), PatchTST (Nie et al. 2023), Crossformer (Wang et al. 2022), Graph Wavenet (Wu et al. 2019), MTGNN (Wu et al. 2020), StemGNN (Cao et al. 2020), CrossGNN (Huang et al. 2023), and FourierGNN (Yi et al. 2023). (2) IMTS classification/imputation models: including GRU-D (Che et al. 2018), SeFT (Horn et al. 2020), RainDrop (Zhang et al. 2022), Warpformer (Zhang et al. 2023), and mTAND (Shukla and Marlin 2021). (3) IMTSF models: including Latent ODEs (Rubanova, Chen, and Duvenaud 2019), CRU (Schirmer et al. 2022), and Neural Flows (Bilos et al. 2021), T-PATCHGNN (Zhang et al. 2024), and TimeCHEAT (Liu, Cao, and Chen 2025). For detailed information on the baseline models, please refer to Appendix A.2.
Implementation Details: We implement APN and all related experiments using the PyTorch framework on a server equipped with an Intel(R) Xeon(R) Gold 6326 CPU (@ 2.90GHz) and an NVIDIA GeForce RTX 3090 GPU. For our APN model, we set the hidden dimension $D$ according to different datasets. Specifically, the hidden dimension $D$ is set to 64 for the PhysioNet and USHCN, while it is set to 16 for the Human Activity and MIMIC. The time encoding dimension $D_{te}$ is uniformly set to 10. The number of patches $P$ is adjusted based on the characteristics of the datasets: 20 for PhysioNet, 16 for MIMIC, 75 for Human Activity, and 4 for USHCN. The training batch size is set to 256 for all datasets. The model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-2}$. We train the model for a maximum of 200 epochs and employ an early stopping strategy with a patience of 50. To ensure the robustness of the results, we repeat each experiment 5 times using five different random seeds. We then report the mean and standard deviation of the performance metrics across these runs.
Dataset PhysioNet MIMIC Human Activity USHCN
[…]
DLinear 41.86 $\pm$ 0.05 15.52 $\pm$ 0.03 4.90 $\pm$ 0.00 16.29 $\pm$ 0.05 4.03 $\pm$ 0.01 4.21 $\pm$ 0.01 6.21 $\pm$ 0.00 3.88 $\pm$ 0.02
TimesNet 16.48 $\pm$ 0.11 6.14 $\pm$ 0.03 5.88 $\pm$ 0.08 13.62 $\pm$ 0.07 3.12 $\pm$ 0.01 3.56 $\pm$ 0.02 5.58 $\pm$ 0.05 3.60 $\pm$ 0.04
PatchTST 12.00 $\pm$ 0.23 6.02 $\pm$ 0.14 3.78 $\pm$ 0.03 12.43 $\pm$ 0.10 4.29 $\pm$ 0.14 4.80 $\pm$ 0.09 5.75 $\pm$ 0.01 3.57 $\pm$ 0.02
Crossformer 6.66 $\pm$ 0.11 4.81 $\pm$ 0.11 2.65 $\pm$ 0.10 9.56 $\pm$ 0.29 4.29 $\pm$ 0.20 4.89 $\pm$ 0.17 5.25 $\pm$ 0.04 3.27 $\pm$ 0.09
Graph Wavenet 6.04 $\pm$ 0.28 4.41 $\pm$ 0.11 2.93 $\pm$ 0.09 10.50 $\pm$ 0.15 2.89 $\pm$ 0.03 3.40 $\pm$ 0.05 5.29 $\pm$ 0.04 3.16 $\pm$ 0.09
MTGNN 6.26 $\pm$ 0.18 4.46 $\pm$ 0.07 2.71 $\pm$ 0.23 9.55 $\pm$ 0.65 3.03 $\pm$ 0.03 3.53 $\pm$ 0.03 5.39 $\pm$ 0.05 3.34 $\pm$ 0.02
StemGNN 6.86 $\pm$ 0.28 4.76 $\pm$ 0.19 1.73 $\pm$ 0.02 7.71 $\pm$ 0.11 8.81 $\pm$ 0.37 6.90 $\pm$ 0.02 5.75 $\pm$ 0.09 3.40 $\pm$ 0.09
CrossGNN 7.22 $\pm$ 0.36 4.96 $\pm$ 0.12 2.95 $\pm$ 0.16 10.82 $\pm$ 0.21 3.03 $\pm$ 0.10 3.48 $\pm$ 0.08 5.66 $\pm$ 0.04 3.53 $\pm$ 0.05
FourierGNN 6.84 $\pm$ 0.35 4.65 $\pm$ 0.12 2.55 $\pm$ 0.03 10.22 $\pm$ 0.08 2.99 $\pm$ 0.02 3.42 $\pm$ 0.02 5.82 $\pm$ 0.06 3.62 $\pm$ 0.07
GRU-D 5.59 $\pm$ 0.09 4.08 $\pm$ 0.05 1.76 $\pm$ 0.03 7.53 $\pm$ 0.09 2.94 $\pm$ 0.05 3.53 $\pm$ 0.06 5.54 $\pm$ 0.38 3.40 $\pm$ 0.28
SeFT 9.22 $\pm$ 0.18 5.40 $\pm$ 0.08 1.87 $\pm$ 0.01 7.84 $\pm$ 0.08 12.20 $\pm$ 0.17 8.43 $\pm$ 0.07 5.80 $\pm$ 0.19 3.70 $\pm$ 0.11
RainDrop 9.82 $\pm$ 0.08 5.57 $\pm$ 0.06 1.99 $\pm$ 0.03 8.27 $\pm$ 0.07 14.92 $\pm$ 0.14 9.45 $\pm$ 0.05 5.78 $\pm$ 0.22 3.67 $\pm$ 0.17
Warpformer 5.94 $\pm$ 0.35 4.21 $\pm$ 0.12 1.73 $\pm$ 0.04 7.58 $\pm$ 0.13 2.79 $\pm$ 0.04 3.39 $\pm$ 0.03 5.25 $\pm$ 0.05 3.23 $\pm$ 0.05
mTAND 6.23 $\pm$ 0.24 4.51 $\pm$ 0.17 1.85 $\pm$ 0.06 7.73 $\pm$ 0.13 3.22 $\pm$ 0.07 3.81 $\pm$ 0.07 5.33 $\pm$ 0.05 3.26 $\pm$ 0.10
Latent-ODE 6.05 $\pm$ 0.57 4.23 $\pm$ 0.26 1.89 $\pm$ 0.19 8.11 $\pm$ 0.52 3.34 $\pm$ 0.11 3.94 $\pm$ 0.12 5.62 $\pm$ 0.03 3.60 $\pm$ 0.12
CRU 8.56 $\pm$ 0.26 5.16 $\pm$ 0.09 1.97 $\pm$ 0.02 7.93 $\pm$ 0.19 6.97 $\pm$ 0.78 6.30 $\pm$ 0.47 6.09 $\pm$ 0.17 3.54 $\pm$ 0.18
Neural Flow 7.20 $\pm$ 0.07 4.67 $\pm$ 0.04 1.87 $\pm$ 0.05 8.03 $\pm$ 0.19 4.05 $\pm$ 0.13 4.46 $\pm$ 0.09 5.35 $\pm$ 0.05 3.25 $\pm$ 0.05
T-PATCHGNN 4.98 $\pm$ 0.08 3.72 $\pm$ 0.03 1.69 $\pm$ 0.03 7.22 $\pm$ 0.09 2.66 $\pm$ 0.03 3.15 $\pm$ 0.02 5.00 $\pm$ 0.04 3.08 $\pm$ 0.04
TimeCHEAT 5.03 $\pm$ 0.07 3.88 $\pm$ 0.03 1.71 $\pm$ 0.02 7.41 $\pm$ 0.10 4.05 $\pm$ 0.10 4.64 $\pm$ 0.07 4.73 $\pm$ 0.07 3.18 $\pm$ 0.05
APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.71 $\pm$ 0.00 3.10 $\pm$ 0.01 4.52 $\pm$ 0.06 2.92 $\pm$ 0.05
Table 1: Overall performance evaluated by MSE and MAE (mean $\pm$ std). The best-performing and second-best results are highlighted in bold and underline, respectively.
@@ 4.2 Main results
We compare the APN model with nineteen baseline models on four challenging datasets—see Table 1. We have the following observations: 1) APN achieves leading prediction accuracy across the board. On all datasets, APN achieves optimal or highly competitive forecasting performance. Compared to the second-best performing model, T-PATCHGNN, APN achieves significant reductions in MSE and MAE metrics by approximately 9.63% and 5.24%, respectively. 2) APN demonstrates exceptional cross-domain generalization capability and robustness. Whether on IMTS datasets with different characteristics in healthcare (PhysioNet, MIMIC), biomechanics (Human Activity), or climate science (USHCN), APN consistently performs excellently. The outstanding performance of APN can be attributed to its TAPA module, which generates superior patch representations by directly performing a soft-aggregation over raw observations within adaptively learned temporal windows. This approach effectively captures salient local dynamics without the potential distortion from interpolation. Building on these high-quality representations, the streamlined query module and MLP decoder ensure a lightweight yet powerful framework.
4.3 Ablation Studies
We perform ablation studies to isolate and validate the contribution of each primary component of APN. The complete results are provided in the Appendix B.2, with a representative summary presented in Table 2. The experiments yield three key observations. First, removing the Adaptive Patching mechanism and reverting to fixed-size windows results in a consistent decline in performance, which highlights the value of dynamically learned boundaries. Second, ablating the Weighted Aggregation scheme for a simple average over hard-cut patches leads to the most significant performance degradation. This empirically confirms that our soft-aggregation design is critical for preventing information loss at patch boundaries (i.e., the coverage issue). Third, replacing the Query-Based Patch Aggregation with a basic linear layer also diminishes performance, proving that our sophisticated aggregation method is essential for integrating contextual information within variables.
Dataset PhysioNet MIMIC USHCN
[…]
@@ 4.3 Ablation Studies
Patch Aggregation
9.92 $\pm$ 0.10 3.57 $\pm$ 0.03 5.21 $\pm$ 0.04
APN (ours) 4.48 $\pm$ 0.06 1.61 $\pm$ 0.02 4.52 $\pm$ 0.06
Table 2: Ablation studies for APN in terms of the best result highlighted in bold.
@@ 4.4 Parameter Sensitivity
Figure 3: Parameter sensitivity studies of main hyper-parameters in APN
Figure 4: We evaluate the memory usage (GB), number of parameters (M), average training time per epoch (s), and total inference time (s) of APN and three representative baselines for IMTSF. All experiments are conducted on the PhysioNet dataset with a consistent batch size of 32 to ensure fair comparison. A lower value indicates better performance.
To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Figure 3. We have the following observations: 1) From Figures 3(a) and 3(b), we observe that for most datasets, a larger $P$ does not necessarily lead to better performance. This suggests that an appropriate $P$ value can better balance the capture of local information and model complexity. 2) Figure 3(c) reveals the impact of the model’s hidden dimension ($D$). In general, a smaller hidden dimension may already be sufficient to capture effective information, while an excessively large dimension can sometimes lead to a slight performance decline. This indicates that the choice of $D$ should be comprehensively considered based on the scale and sparsity of the dataset. 3) Figure 3(d) shows the influence of the time encoding dimension ($D_{te}$). The results indicate that a moderate $D_{te}$ value can provide the model with effective temporal information. An excessively high time encoding dimension may introduce too much noise, interfering with the learning of core features. 4) The APN model exhibits a certain sensitivity to the selection of these core hyperparameters, but it generally achieves competitive results within a reasonable range. In practical applications, appropriate tuning should be performed based on the specific characteristics of the dataset.

### v4

@@ 3.4 Query-based Aggregation
$s_{p}^{n}$ $=\frac{q^{n}\cdot(h_{pe,p}^{n})^{T}}{\sqrt{D}}\quad\text{for }p=1,\dots,P$ (7)
$\beta^{n}$ $=\text{Softmax}(\{s_{p}^{n}\}_{p=1}^{P})$ (8)
$H_{c}^{n}$ $=\sum_{p=1}^{P}\beta_{p}^{n}h_{pe,p}^{n}$ (9)
The resulting summary representations for all channels, $H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{c}\in\mathbb{R}^{N\times D}$. A layer normalization is then applied to stabilize the input for the decoder. This lightweight aggregation mechanism is a direct testament to our design philosophy: by front-loading the complexity into TAPA, the subsequent modules can be elegantly simple yet powerful.
@@ 3.5 Forecasting Decoder
The final step of the APN framework is the ultimate demonstration of its architectural elegance. Having distilled the entire irregular history into a powerful, fixed-size representation $H_{c}^{n}$, the forecasting task becomes remarkably straightforward.
The prediction for a query time $q_{k}^{n}$ is made by a simple two-layer MLP decoder. This decoder takes the concatenated summary representation $H_{c}^{n}$ and the learnable temporal encoding of the query time, $TE(q_{k}^{n})$, as input to produce the final value $\hat{v}^{n}_{k}$:
$\hat{v}^{n}_{k}=\text{MLP}(\text{Concat}(H_{c}^{n},TE(q_{k}^{n})))\in\mathbb{R}$ (10)
@@
$\mathcal{L}=\frac{1}{\sum_{n=1}^{N}Q_{n}}\sum_{n=1}^{N}\sum_{k=1}^{Q_{n}}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (11)
where $Q_{n}$ is the number of query points for each channel.
@@ 4 Experiments
Dataset # Variables # samples Avg # Obs. Max Length
PhysioNet 36 11,981 308.6 47
MIMIC 96 21,250 144.6 96
HumanActivity 12 1,359 362.2 131
USHCN 5 1,114 313.5 337
Table 1: Dataset statistics.
Dataset HumanActivity USHCN PhysioNet MIMIC
Metric MSE MAE MSE MAE MSE MAE MSE MAE
PrimeNet 4.2507±0.0041 1.7018±0.0011 0.4930±0.0015 0.4954±0.0018 0.7953±0.0000 0.6859±0.0001 0.9073±0.0001 0.6614±0.0001
NeuralFlows 0.1722±0.0090 0.3150±0.0094 0.2087±0.0258 0.3157±0.0187 0.4056±0.0033 0.4466±0.0027 0.6085±0.0101 0.5306±0.0066
CRU 0.1387±0.0073 0.2607±0.0092 0.2168±0.0162 0.3180±0.0248 0.6179±0.0045 0.5778±0.0031 0.5895±0.0092 0.5151±0.0048
mTAN 0.0993±0.0026 0.2219±0.0047 0.5561±0.2020 0.5015±0.0968 0.3809±0.0043 0.4291±0.0035 0.9408±0.1126 0.6755±0.0459
SeFT 1.3786±0.0024 0.9762±0.0007 0.3345±0.0022 0.4083±0.0084 0.7721±0.0021 0.6760±0.0029 0.9230±0.0015 0.6628±0.0008
GNeuralFlow 0.3936±0.1585 0.4541±0.0841 0.2205±0.0421 0.3286±0.0412 0.8207±0.0310 0.6759±0.0100 0.8957±0.0209 0.6450±0.0072
GRU-D 0.1893±0.0627 0.3253±0.0485 0.2097±0.0493 0.3045±0.0305 0.3419±0.0029 0.3992±0.0011 0.4759±0.0100 0.4526±0.0055
Raindrop 0.0916±0.0072 0.2114±0.0072 0.2035±0.0336 0.3029±0.0264 0.3478±0.0019 0.4044±0.0020 0.6754±0.1829 0.5444±0.0868
Warpformer 0.0449±0.0010 0.1228±0.0018 0.1888±0.0598 0.2939±0.0591 0.3056±0.0011 0.3661±0.0016 0.4302±0.0035 0.4025±0.0014
tPatchGNN 0.0443±0.0009 0.1247±0.0031 0.1885±0.0403 0.3084±0.0479 0.3133±0.0053 0.3697±0.0049 0.4431±0.0115 0.4077±0.0088
GraFITi 0.0437±0.0005 0.1221±0.0017 0.1691±0.0093 0.2777±0.0248 0.3075±0.0015 0.3637±0.0036 0.4359±0.0455 0.4142±0.0297
APN (Ours) 0.0421±0.0001 0.1159±0.0006 0.1590±0.0137 0.2611±0.0167 0.3093±0.0011 0.3650±0.0026 0.4292±0.0027 0.4016±0.0016
Table 2: Forecasting performance on four IMTS datasets. Overall performance is evaluated by MSE and MAE (mean ± std). The best and second-best results are highlighted in bold and with an underline, respectively.
@@ 4.1 Setup
Datasets: To evaluate the model’s performance, we select four widely used IMTS datasets, including PhysioNet, MIMIC, HumanActivity, and USHCN. These datasets span multiple domains such as healthcare, biomechanics, and climate science. Table 1 summarizes the key statistical features of these datasets. The PhysioNet Challenge 2012 dataset provides clinical time series from the first 48 hours of ICU stays. MIMIC is a large database containing de-identified health data from ICU patients. The HumanActivity dataset consists of biomechanics data with 3D positional variables captured from subjects performing various activities. For climate science, the USHCN dataset includes historical meteorological data from stations across the United States. All datasets are partitioned into training, validation, and test sets using a standard 80%, 10%, and 10% ratio, respectively.
Baselines: To evaluate the performance of APN, we compare it with eleven baseline models. These baseline models can be broadly categorized into two groups: (1) IMTS Classification/Imputation Models: including PrimeNet (PrimeNet), SeFT (SeFT), mTAN (mTAN), GRU-D (GRU_D), Raindrop (RainDrop), and Warpformer (Warpformer). (2) IMTS Forecasting Models: including NeuralFlows (NeuralFlows), CRU (CRU), GNeuralFlow (GNeuralFlow), tPatchGNN (tPatchGNN), GraFITi (GraFITi).
Implementation Details: All our experiments were conducted on a server equipped with an NVIDIA A800 GPU and implemented using the PyTorch 2.6.0+cu124 framework. All models are trained using the Mean Squared Error (MSE) as the loss function and optimized with the AdamW optimizer. We set the maximum number of training epochs to 200 and employ an early stopping strategy, where training is terminated if the model’s performance on the validation set does not improve for 10 consecutive epochs. To ensure a fair comparison, we primarily adopted the hyperparameter settings reported in the original papers for the baseline models. Building on these configurations, we conducted a further comprehensive search and fine-tuning of key hyperparameters on the validation set for all models, including our proposed APN, to ensure each achieved a competitive level of performance. To ensure reproducibility and mitigate the effects of randomness, each experiment is run independently with five different random seeds (from 2024 to 2028), and we report the mean and standard deviation. Detailed hyperparameter configurations for all models are provided in our code repository. We do not apply the “Drop Last” trick (qiu2024tfb; qiu2025tab) to ensure a fair comparison.
@@ 4.2 Main results
We compare the APN model with eleven baseline models on four challenging datasets—see Table 2. We have the following observations: 1) APN achieves leading prediction accuracy across the board. On all datasets, APN achieves optimal or highly competitive forecasting performance. Compared to the second-best performing model, GraFITi, APN achieves significant reductions in MSE and MAE metrics by approximately 2.64% and 3.61%, respectively. 2) APN demonstrates exceptional cross-domain generalization capability and robustness. Whether on IMTS datasets with different characteristics in healthcare (PhysioNet, MIMIC), biomechanics (HumanActivity), or climate science (USHCN), APN consistently performs excellently. The outstanding performance of APN can be attributed to its TAPA module, which generates superior patch representations by directly performing a soft-aggregation over raw observations within adaptively learned temporal windows. This approach effectively captures salient local dynamics without the potential distortion from interpolation. Building on these high-quality representations, the streamlined query module and MLP decoder ensure a lightweight yet powerful framework.
Method MIMIC USHCN
APN (Ours) 0.4292 $\pm$ 0.0027 0.1590 $\pm$ 0.0137
@@
Aggregation
0.4672 $\pm$ 0.0051 0.1981 $\pm$ 0.0327
Table 3: Ablation studies for APN (MSE). Results are reported as mean $\pm$ std. Best results are highlighted in bold.
4.3 Ablation Studies
We perform ablation studies to isolate and validate the contribution of each primary component of APN. The experiments, summarized in Table 3, yield three key observations. First, replacing the Query-Based Aggregation with a basic linear layer leads to the most significant performance degradation. This demonstrates that a simple summarization is insufficient and that our sophisticated query mechanism is essential for dynamically identifying and prioritizing the most salient historical patterns for the forecasting task. Second, ablating the Weighted Aggregation scheme and reverting to a simple average over hard-cut patches also results in a substantial decline in performance. This empirically confirms that our soft-aggregation design is critical for preventing information loss at patch boundaries and ensuring complete information coverage. Finally, removing the Adaptive Patching mechanism in favor of fixed, randomly partitioned windows consistently diminishes performance, highlighting the value of learning dynamic, data-driven temporal boundaries over imposing rigid, arbitrary ones.
Figure 3: Parameter sensitivity studies of main hyperparameters in APN.
Figure 4: Comparison of computational efficiency for APN and three representative baselines. We evaluate four key metrics: (a) peak GPU memory (GB) during a single training step, (b) total number of parameters (M), (c) average training time per step (ms), and (d) average inference time per step (ms). All experiments are conducted on the USHCN dataset with a consistent batch size of 32 to ensure a fair comparison. For all metrics, a lower value indicates better performance.
@@ 4.4 Parameter Sensitivity
To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Figure 3. We have the following observations: 1) Figure 3(c) reveals the impact of the model’s hidden dimension ($D$). For most datasets, performance remains relatively stable across different dimensions, indicating that an excessively large dimension does not bring significant improvement. This suggests that a moderately sized $D$ is sufficient to capture effective information without unnecessary complexity. 2) Figure 3(d) shows the influence of the time encoding dimension ($D_{te}$). The model exhibits greater sensitivity to this parameter, as performance on datasets like PhysioNet and MIMIC degrades significantly when $D_{te}$ becomes too large. This indicates that a moderate $D_{te}$ is optimal for providing effective temporal information, while a higher dimension may introduce noise. 3) From Figures 3(a) and 3(b), we observe the effect of the patch number ($P$). For most datasets, model performance is not highly sensitive to the number of patches, suggesting that the model can robustly capture local features across a reasonable range of $P$ values. 4) The APN model exhibits a certain sensitivity to these core hyperparameters, especially $D_{te}$, but it generally achieves competitive results within a reasonable range. In practical applications, appropriate tuning should be performed based on the specific characteristics of the dataset.